Nightly Observing Plan for MJD 56618 (starting evening of Nov 21, 2013)¶
|Observer 1||Observer 2||Run Manager||SISPI Expert on Call|
|Arnau Pujol||Kevin Reil||Marisa March|
|ext. 511||ext. 513||ext. 516 or 630 338 9267|
CTIO staff (Telescope Operators)¶
|Day Support||Night Assistant|
|Manuel Hernandez||Alberto Alvarez|
DECam Configuration and Setup¶
|SISPI Configuration File||DES.ini|
|Image Processing (Calibration)||NCSA|
|Image Processing (First Cut)||NCSA|
mjd 56618, date 20131121, Site CTIO, UT-offset 3 hr
Current local time 13:57 CLST UT 16:57 SidTim 16:19
Sunset 20:30 CLST 23:30 UT SidTim 343 deg
Eve. -10 21:09 CLST 00:09 UT SidTim 353 deg Zeropoint; standards
Eve. -14 21:31 CLST 00:31 UT SidTim 358 deg Begin obstac
Eve. -18 21:54 CLST 00:54 UT SidTim 4 deg Last chance standards
Moonrise 23:41 CLST 02:41 UT SidTim 30 deg
Midpoint 01:29 CLST 04:29 UT SidTim 58 deg
Morn -18 05:04 CLST 08:04 UT SidTim 111 deg Begin astronomical twilight
Morn Std 05:26 CLST 08:26 UT SidTim 117 deg End obstac; begin standards
Morn -10 05:48 CLST 08:48 UT SidTim 122 deg
Sunrise 06:28 CLST 09:28 UT SidTim 132 deg
Moon position (deg): RA 111.9, Dec 17.5 Phase: 0.711
Clear, some cloud on horizon.
Observing Plan and Goals for the Night¶
- Afternoon calibrations script to use: prenight_calibration.json
- Start of night standards.
- Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
- End of night standards.
- End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json
Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.
UT Date = '2013/11/22 00:09' Local time = 21:09
E2-A ra 04:03:00 dec -44:41:45 secz 1.7 script to use: DES_std_e2a_Yzirg
SDSSJ0200-0000 ra 02:00:00 dec +00:00:00 secz 1.5 script to use: DES_std_sdssj0200_Yzirg
SDSSJ2300-0000 ra 23:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj2300_Yzirg
End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/11/22 08:26' Local time = 05:26
C26202/HST ra 03:32:30 dec -27:46:05 secz 1.8 script to use: DES_std_c26202_grizY
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.6 script to use: DES_std_sdssj1048_grizY
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_grizY
Notes from 4pm Meeting¶
(Thanks to Tom for minutes)
Run Manager's Meeting Nov. 21, 2013
Tom, Marisa, Robert, Eric, Manuel, Don, Margaret, Alex
Camera and Telescope
Telescope is on white spot and camera is ready to go.
Last Night's Data
WF and SN the night before.
Last night was very nice. WF z and Y-band with < 1" seeing.
Nightly out-of-focus images are working and will be done regularly.
Last night had to restart pipeline6, observer1, and readout1 machines
Got last night's data and it is being processed already.
DQ is in hand for 11/18 and 11/19.
X1 from a couple nights ago (11/18) data quality failed seeing
and signal to noise for 2 of the filters => X1 needs to be redone.
Robert's version of the software flagged that as so. Robert produced
the DQ file and sent it to DES operations for implementation.
OBSTAC decides what to observe, including whether to do SN sequence.
SN would be happy to allow one image of a SN sequence to be bad but
still keep the sequence as good, but at present OBSTAC will flag the sequence
as bad if any of the images is labeled as bad. In tonight's case, since
2 filters of X1 were bad, everybody is happy with the result. E.N. will
work on OBSTAC in this regard but that will take some time.
Data bases going down starting on Friday afternoon. Will be able to
receive data but not process it during the interval.
Weather looks good.
We expect it to redo X1.
Otherwise g,r,i when the moon is down, i,z,Y when the Moon comes up.